# Leveraging (Biased) Information: Multi-armed Bandits with Offline Data

### v1

@@ 4 Design and Analysis of MIN-UCB
After the impossibility result, we present our MIN-UCB policy in Algorithm 1. MIN-UCB follows the optimism-in-face-of-uncertainty principle. At time $t>K$, MIN-UCB sets $\min\{\text{UCB}_{t}(a),\text{UCB}_{t}^{\text{S}}(a)\}$ as the UCB on $\mu^{\text{(on)}}(a)$. While $\text{UCB}_{t}(a)$ follows (Auer et al., 2002), the construction of $\text{UCB}_{t}^{\text{S}}(a)$ in (6) adheres to the following ideas that incorporates a valid bias bound $V$, the auxiliary input.
@@
6:     For all $a\in\mathcal{A}$, compute the warm-start $\text{UCB}^{\text{S}}_{t}(a)=$
@@
7:     Select
$A_{t}\in\text{argmax}_{a\in{\cal A}}\left\{\min\left\{\text{UCB}_{t}(a),\text{UCB}^{\text{S}}_{t}(a)\right\}\right\}.$
@@
Finally, we adopt $\text{UCB}_{t}^{\text{S}}(a)$ as the UCB when the bias bound $V(a)$ is so small and the sample size $T_{\text{S}}(a)$ is so large that $\text{UCB}_{t}(a)>\text{UCB}_{t}^{\text{S}}(a)$. Otherwise, $P^{\text{(off)}}$ is deemed “far away” $P^{\text{(on)}}$, and we follow the vanilla UCB. Our adaptive decisions on whether to incorporate offline data at each time step is different from the vanilla UCB policy that always ignores offline data, or from HUCB and MonUCB that always incorporate all offline data.
Lastly, we remark on $V$. Suppose that the DM possesses no knowledge on $P^{\text{(off)}},P^{\text{(on)}}$ additional to the offline dataset $S$ at the start of the online phase. Then, it is advisable to set $V(a)=\infty$ for all $a$, which specializes MIN-UCB to the vanilla UCB policy. Indeed, the theoretical impossibility result in Section 3 dashes out any hope for adaptively tuning $V$ in MIN-UCB with online data, in order to outperform the vanilla UCB. By contrast, when the DM has a non-trivial upper bound $V(a)$ on $\left|\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)\right|$, we show that can uniformly out-perform the vanilla UCB that direclty ignores the offline data.
We next embark on the analysis on MIN-UCB, by establishing instance dependent regret bounds in Section 4.1 and instance independent regret bounds in Section 4.3. Both sections relies on considering the following events of accurate estimations by $\text{UCB}_{t}(a),\text{UCB}^{\text{S}}_{t}(a)$. For each $t$, define ${\cal E}_{t}=\cap_{a\in{\cal A}}({\cal E}_{t}(a)\cap{\cal E}^{\text{S}}_{t}(a))$, where
${\cal E}_{t}(a)=\left\{\mu^{\text{(on)}}(a)\leq\text{UCB}_{t}(a)\leq\mu^{\text{(on)}}(a)+2\text{rad}_{t}(a)\right\},$
${\cal E}^{\text{S}}_{t}(a)=\left\{\mu^{\text{(on)}}(a)\leq\text{UCB}^{\text{S}}_{t}(a)\leq\mu^{(\text{on})}(a)+\text{rad}^{\text{S}}_{t}(a)\right.$
$\left.+\left[\sqrt{\frac{2\log(2t/\delta_{t})}{N_{t}(a)+T_{\text{S}}(a)}}+\frac{T_{\text{S}}(a)\cdot(\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a))}{N_{t}(a)+T_{\text{S}}(a)}\right]\right\}.$
While ${\cal E}_{t}(a)$ incorporates estimation error due to stochastic variations, ${\cal E}^{\text{S}}_{t}(a)$ additionally involves the potential bias of the offline data.
Lemma 4.1.
$\Pr({\cal E}_{t})\geq 1-2K\delta_{t}.$
The proof of Lemma 4.1 is provided in Appendix B.1.
4.1 Instance Dependent Regret Bounds
Both our instance dependent regret upper and lower bounds depend on the following discrepancy measure
@@ 4.1 Instance Dependent Regret Bounds
on an arm $a$. Note that $\omega(a)$ depends on both the valid bias bound $V(a)$ and the model parameters $\mu^{\text{(off)}}(a),\mu^{\text{(on)}}(a)).$ By the validity of $V$, we know that $\omega(a)\in[0,2V(a)]$.
Theorem 4.2.
Consider policy MIN-UCB, which inputs a valid bias bound $V$ on the underlying instance $I$ and $\delta_{t}=\frac{1}{2Kt^{2}}$ for $t=1,2,\ldots$. We have $\mathbb{E}[\text{Reg}_{T}(\text{MIN-UCB},P)]\leq$
@@
Theorem 4.2 is proved in the Appendix B.2. We provide a sketch proof in Section 4.2. A full explicit bound is in (28) in the Appendix. Comparing with the regret bound (1) by the vanilla UCB, MIN-UCB provides an improvement on the regret order bound by the saving term $\text{Sav}_{0}(a)$.
First, note that $\text{Sav}_{0}(a)\geq 0$ always. The case of $\omega(a)<\Delta(a)$ means that the reward distributions $P^{\text{(off)}}_{a},P^{\text{(on)}}_{a}$ are “sufficiently close”. The offline data on arm $a$ are incorporated by MIN-UCB to improve upon the vanila UCB. By contrast, the case of $\omega(a)\geq\Delta(a)$ means that $P^{\text{(off)}}_{a},P^{\text{(on)}}_{a}$ are “far apart”, so that the offline data on arm $a$ are ignored in the learning process, while MIN-UCB still matches the performance guarantee of the vanilla UCB. We emphasize that $\omega(a)$ and $\Delta(a)$ are both latent, so MIN-UCB conducts the above adaptive choice under uncertainty.
Next, in the case of $\omega(a)<\Delta(a)$ where the offline data on $a$ are beneficial, observe that $\text{Sav}_{0}(a)$ is increasing in $T_{\text{S}}(a)$. This monotonicity adheres to the intuition that more offline data leads to a higher reward when $P^{\text{(off)}}_{a},P^{\text{(on)}}_{a}$ are “sufficiently close”. The term $\text{Sav}_{0}(a)$ dcreases as $\omega(a)$ increses. Indeed, a smaller $V(a)$ means a tighter upper bound on $|\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)|$ is known to the DM, which leads to a better peformance.
Interestingly, $\text{Sav}_{0}(a)$ increases when $\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)$ decreases. Paradoxically, $\text{Sav}_{0}(a)$ increases with $|\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)|$, under the case of $\mu^{\text{(off)}}(a)<\mu^{\text{(on)}}(a)$. Indeed, the saving term $\text{Sav}_{0}(a)$ depends not only on the magnitude of the distributional shift, but also the direction. To gain intuition, consider the case of $\mu^{\text{(off)}}(a)<\mu^{\text{(on)}}(a)$. On $\mu^{\text{(on)}}(a)$, the offline dataset on $a$ provides a pessimistic estimate, which encourages the DM to eliminate arm $a$ (which is sub-optimal), hence improves the DM’s performance.
Lastly, we compare to existing baselines. By setting $V(a)=\infty$ for all $a$, the regret bound (9) reduces to (1) of the vanilla UCB. In the case when $P^{\text{(on)}}=P^{(\text{off})}$ and it is known to the DM, setting $V(a)=0$ for all $a$ reduces the regret bound (9) to (2). In both cases we achieve the state-of-the-art regret bounds, while additionally we quantify the impact of distributional drift to the regret bound. Finally, the ARROW-CB by (Zhang et al., 2019) can be applied to our model. While ARROW-CB does not require any auxiliary input, their regret bound is at least $\Omega(K^{1/3}T^{2/3})$. ARROW-CB provides improvment to purely online policies in the contextual bandit setting, as discussed in Appednix A.1.
@@
Knowing a valid bias bound $V$ is equivalent to knowing that the underlying instance $I$ lies in the subset ${\cal I}_{V}$, leading to the following which is equivalent to Theorem 4.2:
Corollary 4.4.
@@
Definition 4.6.
For $C>0,p\in(0,1)$ and a collection ${\cal I}$ of instances, a non-anticipatory policy $\pi$ is said to be $(C,p)$-consistent on ${\cal I}$, if for all $I\in{\cal I}$, it holds that $\mathbb{E}[\text{Reg}_{T}(P,\pi)]\leq CT^{p}$.
For example, the vanilla UCB is $(C,1/2)$ consistent for some absolute constant $C>0$.
Theorem 4.7.
@@
holds for any $\epsilon\in(0,1]$, where
$\text{Sav}_{\epsilon}(a)=T_{\text{S}}(a)\cdot\Delta(a)\cdot\max\left\{1-\frac{\omega(a)}{(1+\epsilon)\Delta(a)},0\right\}^{2},$ (12)
and $\kappa_{\epsilon}(a)=\frac{1}{2(1+\epsilon)^{2}\Delta(a)}\log\frac{\epsilon\Delta(a)}{8C}$ is independent of $T$ or $T_{\text{S}}(a)$.
Theorem 4.7 is proved in Appendix C.3. Modulo the additive term $\kappa_{\epsilon}(a)$, the regret upper bound (9) and lower bound (11) are nearly matching. Both bounds feature the $\Theta((\log T)/\Delta(a))$ regret term that is classical in the literature (Lattimore & Szepesvári, 2020). Importantly, the lower bound (11) features the regret saving term $\text{Sav}_{\epsilon}(a)$, which closely matches the term $\text{Sav}_{0}(a)$ in the upper bound. Indeed, $\epsilon\in(0,1]$ is arbitrary, $\text{Sav}_{\epsilon}(a)$ tends to the term $\text{Sav}_{0}(a)$ in the regret upper bound (9) as we tend $\epsilon$ to 0. The close match highlights the fundamental nature of the discrepancy measure $\omega(a)$. In addition, the lower bound (11) carries the insight that the offline data on arm $a$ facilitates improvement over the vanilla UCB if and only if $\omega(a)<(1+\epsilon)\Delta(a)$, which is consistent with our insights from the regret upper bound (9) by tending $\epsilon\rightarrow 0$.
Lastly, in the case $P^{\text{(off)}}=P^{\text{(on)}}$, the lower bound (11) reduces to $\Omega(\sum_{a:\Delta(a)>0}[\frac{\log T}{\Delta(a)}-T_{\text{S}}(a)\Delta(a)+\kappa_{\epsilon}(a)])$, which ascertains the near tightness of (2) by HUCB1 and MonUCB. A nearly tight regret lower bound is novel in the literature.
4.2 Sketch Proof of Theorem 4.2
Theorem 4.2 is a direct conclusion of the following lemma:
Lemma 4.8.
Let $a$ be a sub-optimal arm, that is $\Delta(a)>0$. Conditioned on the event ${\cal E}_{t}$, if it holds that
$N_{t}(a)>32\cdot\frac{\log(4Kt^{4})}{\Delta(a)^{2}}-T_{\text{S}}(a)\cdot\max\left\{1-\frac{\omega(a)}{\Delta(a)},0\right\}^{2},$
@@
Note $\frac{\text{Sav}_{0}(a)}{\Delta(a)}=T_{\text{S}}(a)\cdot\max\{1-\frac{\omega(a)}{\Delta(a)},0\}^{2}$. The Lemma is fully proved in Appendix B.2, and a sketch proof is below.
Sketch Proof of Lemma 4.8.
Consider
Case 1a: $\omega(a)<\Delta(a)$, and $\frac{\text{Sav}_{0}(a)}{\Delta(a)}\geq\frac{16\log(4Kt^{4})}{\Delta(a)^{2}}$,
Case 1b: $\omega(a)<\Delta(a)$, and $\frac{\text{Sav}_{0}(a)}{\Delta(a)}<\frac{16\log(4Kt^{4})}{\Delta(a)^{2}}$,
Case 2: $\omega(a)\geq\Delta(a)$.
Both Cases 1b, 2 imply $N_{t}(a)>16\cdot\frac{\log(4Kt^{4})}{\Delta(a)^{2}}$. Following (Auer et al., 2002), we can then deduce
$\text{UCB}_{t}(a)$ $\leq\mu^{\text{(on)}}(a)+\frac{\Delta(a)}{2}$
$<\mu^{(\text{on})}_{*}\leq\min\left\{\text{UCB}_{t}(a_{*}),\text{UCB}^{\text{S}}_{t}(a_{*})\right\},$
thus $A_{t}\neq a$ with certainty. In what follows, we focus on Case 1a, the main case when MIN-UCB incorporates the offline dataset in learning. Case 1a and $V(a)$ being a valid bias bound imply
$\frac{\omega(a)}{\Delta(a)}\in[0,1).$ (13)
Then
@@
$\overset{(\dagger)}{\leq}$ $\left(1-\frac{\omega(a)}{\Delta(a)}\right)\sqrt{\frac{2\log(4Kt^{4})}{\frac{16\log(4Kt^{4})}{\Delta(a)^{2}}}}=\left(1-\frac{\omega(a)}{\Delta(a)}\right)\cdot\frac{\Delta(a)}{\sqrt{8}}.$
()($\dagger$)( † ) is by the second condition in Case 1a and (13). Also,
$\frac{T_{\text{S}}(a)\cdot\omega(a)}{N_{t}(a)+T_{\text{S}}(a)}\leq\frac{\omega(a)}{\Delta(a)}\cdot\Delta(a).$
Consequently, conditioned on event ${\cal E}_{t}$,
$\text{UCB}^{\text{S}}_{t}(a)$
$\leq$ $\mu^{\text{(on)}}(a)+2\sqrt{\frac{2\log(2t/\delta_{t})}{N_{t}(a)+T_{\text{S}}(a)}}+\frac{T_{\text{S}}(a)\cdot\omega(a)}{T_{\text{S}}(a)+N_{t}(a)}$ (14)
@@
$\overset{(\ddagger)}{<}$ $\mu^{(\text{on})}_{*}\overset{(\P)}{\leq}\min\left\{\text{UCB}_{t}(a_{*}),\text{UCB}^{\text{S}}_{t}(a_{*})\right\},$
()($\ddagger$)( ‡ ) is because $10\leq 1-\frac{\omega(a)}{\Delta(a)}<$ 1 by (13). ($\P$, 14) are by the event ${\cal E}_{t}$. Altogether, $A_{t}\neq a$ with certainty. ∎
4.3 Instance Indepedent Regret Bounds
Now we analyze the instance independent bound. We set $\delta_{t}=\delta/(2Kt^{2})$, where $\delta\in(0,1)$ is a prescribed confidence paramater.
Theorem 4.9.
Assume $2\leq K\leq T$. Consider policy MIN-UCB, which inputs a valid bias bound $V$ on the underlying instance and $\delta_{t}=\delta/(2Kt^{2})$ for $t=1,\ldots$, where $\delta\in(0,1)$. With probability at least $1-\delta$,  $\text{Reg}_{T}(\text{MIN-UCB},P)=$
$O\left(\min\left\{\sqrt{KT\log\frac{T}{\delta}},\left(\sqrt{\frac{\log\frac{T}{\delta}}{\tau_{*}}}+V_{\text{max}}\right)\cdot T\right\}\right),$ (15)

### v2

@@ 4  MIN-UCB: Leverage Offline Data with Auxiliary Information
We present our MIN-UCB policy in Algorithm 1. MIN-UCB follows the optimism-in-face-of-uncertainty principle. At time $t>K$, MIN-UCB sets $\min\{\text{UCB}_{t}(a),\text{UCB}_{t}^{\text{S}}(a)\}$ as the UCB on $\mu^{\text{(on)}}(a)$. While $\text{UCB}_{t}(a)$ follows AuerCBF02, the construction of $\text{UCB}_{t}^{\text{S}}(a)$ in (8) adheres to the following ideas that incorporates a valid bias bound $V$, the auxiliary input.
@@
6:   For all $a\in\mathcal{A}$, compute the warm-start UCB: $\text{UCB}^{\text{S}}_{t}(a)=$
@@
7:    Select $A_{t}\in\text{argmax}_{a\in{\cal A}}\{\min\{\text{UCB}_{t}(a),\text{UCB}^{\text{S}}_{t}(a)\}\}.$
@@
Finally, we adopt $\text{UCB}_{t}^{\text{S}}(a)$ as the UCB when the bias bound $V(a)$ is so small and the sample size $T_{\text{S}}(a)$ is so large such that $\text{UCB}_{t}(a)>\text{UCB}_{t}^{\text{S}}(a)$. Otherwise, $P^{\text{(off)}}$ is “far away” $P^{\text{(on)}}$, and we follow the vanilla UCB. Our adaptive decisions on whether to incorporate offline data at each time step is different from the vanilla UCB policy that always ignores offline data, or from HUCB and MonUCB that always incorporate all offline data.
We remark on the valid bias bound $V$. Suppose that the DM has no knowledge on $P^{\text{(off)}},P^{\text{(on)}}$ at the start of the online phase. Then, it is advisable to set $V(a)=\infty$ for all $a$, which specializes MIN-UCB to the vanilla UCB. Indeed, the impossibility result in Section 3 dashes out any hope for adaptively tuning $V$ in MIN-UCB with online data, in order to outperform the vanilla UCB. In contrast, when the DM has a non-trivial upper bound $V(a)$ on $\left|\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)\right|$, we show that our algorithm can uniformly outperform the vanilla UCB that direclty ignores the offline data.
We next embark on the analysis on MIN-UCB, by establishing instance-dependent regret bounds in Section 4.1 and instance-independent regret bounds in Section 4.3. Both sections relies on considering the following events of accurate estimations by $\text{UCB}_{t}(a),\text{UCB}^{\text{S}}_{t}(a)$. For each $t$, define $\xi_{t}=\cap_{a\in{\cal A}}(\xi_{t}(a)\cap\xi^{\text{S}}_{t}(a))$, where $\xi_{t}(a)=\left\{\mu^{\text{(on)}}(a)\leq\text{UCB}_{t}(a)\leq\mu^{\text{(on)}}(a)+2\text{rad}_{t}(a)\right\}$, and $\xi^{\text{S}}_{t}(a)=$
$\left\{\mu^{\text{(on)}}(a)\leq\text{UCB}^{\text{S}}_{t}(a)\leq\mu^{(\text{on})}(a)+\text{rad}^{\text{S}}_{t}(a)+\right$.
$\left.\sqrt{\frac{2\log(2t/\delta_{t})}{N_{t}(a)+T_{\text{S}}(a)}}+\frac{T_{\text{S}}(a)\cdot(\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a))}{N_{t}(a)+T_{\text{S}}(a)}\right\}.$
While $\xi_{t}(a)$ incorporates estimation error due to stochastic variations, $\xi^{\text{S}}_{t}(a)$ additionally involves the potential bias of the offline data. The following Lemma 1 is proved in Appendix B.1.
Lemma 1.
$\Pr(\xi_{t})\geq 1-2K\delta_{t}.$
4.1  Instance-Dependent Regret Bounds
Our instance-dependent regret bounds depend on the following discrepancy measure
@@ 4.1  Instance-Dependent Regret Bounds
on an arm $a$. Note that $\omega(a)$ depends on both the valid bias bound $V(a)$ and the model parameters $\mu^{\text{(off)}}(a),\mu^{\text{(on)}}(a).$ By the validity of $V$, we know that $\omega(a)\in[0,2V(a)]$.
Theorem 2.
Consider policy MIN-UCB, which inputs a valid bias bound $V$ on the underlying instance $I$ and $\delta_{t}=1/(2Kt^{2})$ for $t=1,2,\ldots$. We have $\mathbb{E}[\text{Reg}_{T}(\text{MIN-UCB},P)]\leq$
@@
Theorem 2 is proved in the Appendix B.2. We provide a sketch proof in Section 4.2. A full explicit bound is in (32) in the Appendix. Comparing with the regret bound (1) by the vanilla UCB, MIN-UCB provides an improvement on the regret order bound by the saving term $\text{Sav}_{0}(a)$.
First, $\text{Sav}_{0}(a)\geq 0$ always holds. The case of $\omega(a)<\Delta(a)$ means that the reward distributions $P^{\text{(off)}}_{a},P^{\text{(on)}}_{a}$ are “sufficiently close”. The offline data on arm $a$ are incorporated by MIN-UCB to improve upon the vanila UCB. By contrast, the case of $\omega(a)\geq\Delta(a)$ means that $P^{\text{(off)}}_{a},P^{\text{(on)}}_{a}$ are “far apart”, so that the offline data on arm $a$ are ignored in the learning process, while MIN-UCB still matches the performance guarantee of the vanilla UCB. We emphasize that $\omega(a)$ and $\Delta(a)$ are both latent, so MIN-UCB conducts the above adaptive choice under uncertainty.
Next, when $\omega(a)<\Delta(a)$ and the offline data on $a$ are beneficial, note that $\text{Sav}_{0}(a)$ is increasing in $T_{\text{S}}(a)$. This monotonicity adheres to the intuition that more offline data leads to a higher reward when $P^{\text{(off)}}_{a},P^{\text{(on)}}_{a}$ are “sufficiently close”. The term $\text{Sav}_{0}(a)$ decreases as $\omega(a)$ increases. A smaller $V(a)$ means a tighter upper bound on $|\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)|$ is known to the DM, which leads to a better peformance.
Interestingly, $\text{Sav}_{0}(a)$ increases when $\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)$ decreases. Paradoxically, $\text{Sav}_{0}(a)$ increases with $|\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)|$, under the case of $\mu^{\text{(off)}}(a)<\mu^{\text{(on)}}(a)$. Indeed, the saving term $\text{Sav}_{0}(a)$ depends not only on the magnitude of the distributional shift, but also the direction. To gain intuition, consider the case of $\mu^{\text{(off)}}(a)<\mu^{\text{(on)}}(a)$. On $\mu^{\text{(on)}}(a)$, the offline dataset on $a$ provides a pessimistic estimate, which encourages the DM to eliminate the sub-optimal arm $a$, hence improves the DM’s performance.
Lastly, we compare to existing baselines. By setting $V(a)=\infty$ for all $a$, the regret bound (11) reduces to (1) of the vanilla UCB. In the case when $P^{\text{(on)}}=P^{(\text{off})}$ and it is known to the DM, setting $V(a)=0$ for all $a$ reduces the regret bound (11) to (2). In both cases we achieve the state-of-the-art regret bounds, while additionally we quantify the impact of distributional drift to the regret bound. Finally, the ARROW-CB by zhang2019warm can be applied to our model. While ARROW-CB does not require any auxiliary input, their regret bound is at least $\Omega(K^{1/3}T^{2/3})$. ARROW-CB provides improvment to purely online policies in the contextual bandit setting, as discussed in Appednix A.1.
@@
Knowing a valid bias bound $V$ is equivalent to knowing that the underlying instance $I$ lies in the subset ${\cal I}_{V}$, leading to the following which is equivalent to Theorem 2:
Corollary 1.
@@
Definition 3.
[…]
For example, the vanilla UCB is $(C,1/2)$ consistent for some $C$ proportional to $\sqrt{K}$.
Theorem 3.
@@
holds for any $\epsilon\in(0,1]$, where $\text{Sav}_{\epsilon}(a)=$
$T_{\text{S}}(a)\cdot\Delta(a)\cdot\max\left\{1-\frac{\omega(a)}{(1+\epsilon)\Delta(a)},0\right\}^{2},$ (14)
and $\kappa_{\epsilon}(a)=\frac{1}{2(1+\epsilon)^{2}\Delta(a)}\log\frac{\epsilon\Delta(a)}{8C}$ is independent of $T$ and $T_{\text{S}}(a)$.
Theorem 3 is proved in Appendix D.2. Modulo the additive term $\kappa_{\epsilon}(a)$, the regret upper bound (11) and lower bound (13) are nearly matching. Both bounds feature the $\Theta((\log T)/\Delta(a))$ regret term that is classical in the literature lattimore2020bandit. Importantly, the lower bound (13) features the regret saving term $\text{Sav}_{\epsilon}(a)$, which closely matches the term $\text{Sav}_{0}(a)$ in the upper bound. Indeed, $\epsilon\in(0,1]$ is arbitrary, $\text{Sav}_{\epsilon}(a)$ tends to the term $\text{Sav}_{0}(a)$ in the regret upper bound (11) as we tend $\epsilon$ to 0. The close match highlights the fundamental nature of the discrepancy measure $\omega(a)$. In addition, the lower bound (13) carries the insight that the offline data on arm $a$ facilitates improvement over the vanilla UCB if and only if $\omega(a)<(1+\epsilon)\Delta(a)$, which is consistent with our insights from the regret upper bound (11) by tending $\epsilon$ to 0.
Lastly, in the case $P^{\text{(off)}}=P^{\text{(on)}}$, the lower bound (13) reduces to $\Omega(\sum_{a:\Delta(a)>0}[\frac{\log T}{\Delta(a)}-T_{\text{S}}(a)\Delta(a)+\kappa_{\epsilon}(a)])$, which ascertains the near tightness of (2) by HUCB1 and MonUCB. We highlight that a nearly tight regret lower bound in the well-aligned case of $P^{\text{(off)}}=P^{\text{(on)}}$ is novel in the literature.
4.2  Sketch Proof for Theorem 2
Theorem 2 is a direct conclusion of the following lemma:
Lemma 2.
Let $a$ be a sub-optimal arm, that is $\Delta(a)>0$. Conditioned on the event $\xi_{t}$, if it holds that $N_{t}(a)>$
$32\cdot\frac{\log(4Kt^{4})}{\Delta(a)^{2}}-T_{\text{S}}(a)\cdot\max\left\{1-\frac{\omega(a)}{\Delta(a)},0\right\}^{2},$
@@ 4.2  Sketch Proof for Theorem 2
To relate Theorem 2 to Lemma 2, we remark that $\frac{\text{Sav}_{0}(a)}{\Delta(a)}=T_{\text{S}}(a)\cdot\max\{1-\frac{\omega(a)}{\Delta(a)},0\}^{2}$.
Proof of Lemma 2.
We consider two main cases on $\omega(a)$ and $\Delta(a)$:
Case 1: $\omega(a)<\Delta(a)$. In this case, $P^{\text{(off)}}_{a}$ and $P^{\text{(on)}}_{a}$ are sufficiently closed. The case condition implies that $\frac{\omega(a)}{\Delta(a)}\in[0,1)$ and
$\max\left\{1-\frac{\omega(a)}{\Delta(a)},0\right\}^{2}=\left(1-\frac{\omega(a)}{\Delta(a)}\right)^{2}>0.$
To proceed. we consider two sub-cases:
Sub-case 1a: $T_{\text{S}}(a)\cdot(1-\frac{\omega(a)}{\Delta(a)})^{2}\geq\frac{16\log(4Kt^{4})}{\Delta(a)^{2}}$. In this sub-case, $T_{\text{S}}(a)$ is sufficiently large such that the offine data for arm $a$ provides enough information for the algorithm to eliminate this sub-optimal arm. We focus on analyzing $\text{UCB}_{t}^{\text{S}}(a)$:
@@
$\leq\left(1-\frac{\omega(a)}{\Delta(a)}\right)\sqrt{\frac{2\log(4Kt^{4})}{\frac{16\log(4Kt^{4})}{\Delta(a)^{2}}}}$
$=\left(1-\frac{\omega(a)}{\Delta(a)}\right)\cdot\frac{\Delta(a)}{\sqrt{8}}.$
Consequently, conditioned on event $\xi_{t}$, $\text{UCB}^{\text{S}}_{t}(a)\leq\mu^{\text{(on)}}(a)+$
$\qquad\ 2\cdot\sqrt{\frac{2\log(2t/\delta_{t})}{N_{t}(a)+T_{\text{S}}(a)}}+\frac{T_{\text{S}}(a)}{T_{\text{S}}(a)+N_{t}(a)}\cdot\omega(a)$ (15a)
@@
$<\mu^{(\text{on})}_{*}\leq\min\left\{\text{UCB}_{t}(a_{*}),\text{UCB}^{\text{S}}_{t}(a_{*})\right\},$ (15b)
The strict inequality in step (15b) is by the sub-case condition that $1-\frac{\omega(a)}{\Delta(a)}>0$. The less-than-equal steps in Lines (15a, 15b) is by conditioning on the event $\xi_{t}$. Altogether, $A_{t}\neq a$ with certainty.
Sub-case 1b: $T_{\text{S}}(a)\cdot(1-\frac{\omega(a)}{\Delta(a)})^{2}<\frac{16\log(4Kt^{4})}{\Delta(a)^{2}}$. In this sub-case, although $P^{\text{(off)}}_{a}$ and $P^{\text{(on)}}_{a}$ are sufficiently closed, $T_{\text{S}}(a)$ is not large enough for $\text{UCB}_{t}^{\text{S}}(a)$ to shrink faster than $\text{UCB}_{t}(a)$. Thus the offline data does not provide enough evidence for the algorithm to eliminate the sub-optimal arm $a$ more quickly. Then we have $N_{t}(a)>$
[…]
$\geq$ $\frac{16\log(4Kt^{4})}{\Delta(a)^{2}}.$
Consequently, we have
leading to $\text{UCB}_{t}(a)\leq\mu^{\text{(on)}}(a)+2\cdot\frac{\Delta(a)}{\sqrt{8}}<\mu^{(\text{on})}_{*}\leq\min\{\text{UCB}_{t}(a_{*}),\text{UCB}^{\text{S}}_{t}(a_{*})\}$, where the first and the last inequalities are by the event $\xi_{t}$, thus $A_{t}\neq a$ with certainty.
Case 2: $\omega(a)\geq\Delta(a)$. Then $\max\{1-\frac{\omega(a)}{\Delta(a)},0\}^{2}=0$, and $N_{t}(a)>32\cdot\frac{\log(4Kt^{4})}{\Delta(a)^{2}}$. In this case, $P^{\text{(off)}}_{a}$ and $P^{\text{(on)}}_{a}$ are far apart. Similar to Case 1b, we arrive at $\text{rad}_{t}(a)=$
$\sqrt{\frac{2\log(2t/\delta_{t})}{N_{t}(a)}}<\sqrt{\frac{2\log(4Kt^{4})}{32\cdot\frac{\log(4Kt^{4})}{\Delta(a)^{2}}}}\leq\frac{\Delta(a)}{4}\leq\frac{\Delta(a)}{\sqrt{8}}.$
Then same as Case 1b, $\text{UCB}_{t}(a)\leq\min\{\text{UCB}_{t}(a_{*}),\text{UCB}^{\text{S}}_{t}(a_{*})\}$. Altogether, $A_{t}\neq a$ with certainty. Altogether, all cases are covered and the Lemma is proved. ∎
4.3  Instance-Indepedent Regret Bounds
Now we analyze the instance-independent bound.
Theorem 4.
Assume $2\leq K\leq T$. Consider policy MIN-UCB, which inputs a valid bias bound $V$ on the underlying instance and $\delta_{t}=\delta/(2Kt^{2})$ for any $t$, where $\delta\in(0,1)$. With probability at least $1-\delta$,  $\text{Reg}_{T}(\text{MIN-UCB},P)=$
$O\left(\min\left\{\sqrt{KT\log\frac{T}{\delta}},\left(\sqrt{\frac{\log(T/\delta)}{\tau_{*}}}+V_{\text{max}}\right)\cdot T\right\}\right),$ (16)